# Magnetic Bloch oscillations in a non-Hermitian quantum Ising chain

### v1

@@ III.3 Bessel initial state
While the results in Eqs. (27) and (28) indicate that this is a periodic oscillation with period $T=\pi/h_{z}$, the pattern is not so explicit. Nevertheless, for a finite system size $N$, we introduce the center of the wavepacket
@@ IV Numerical simulations
The boundary of $\sigma^{z}_{j}\right>=+1$ and $-1$ is the position of the magnetic domain wall. For the Hermitian spin chain, Figs. 3(a1)-(a3) show that the the dynamics are magnetic Bloch breathing, BO and stationary modes for the Kronecker delta, Gaussian and Bessel initial states, respectively. With the same initial states, for the non-Hermitian spin chain, the results in Figs. 3(b1)-(b3) indicate that the dynamics are magnetic Bloch breathing, stationary and BO modes, respectively. For the later two initial states, the corresponding BOs exhibit totally different patterns for the same initial states in the two different systems. These numerical results are in accordance with the analyses in the previous section.
@@ V Summary
In summary, we demonstrate the existence of the magnetic BOs in a non-Hermitian quantum Ising chain. It is shown that in the small-field region, the low-energy dynamics of the magnetic domain walls are captured by a single-particle effective Hamiltonian, with the transverse field acting as a hopping coefficient for the magnetic domain wall and the strength of the longitudinal field playing the role of a skew potential. For real and imaginary transverse fields, the eigenstates of the effective Hamiltonian are both localized states with equally spaced energy levels, forming the WS ladders. Analytical and numerical calculations of the time evolution for the non-Hermitian quantum Ising chain show that (i) the dynamics of the Kronecker delta initial state follow a breathing mode; While (ii) the Gaussian state remains stationary, which is different from the oscillation of the domain wall in a Hermitian quantum Ising chain; (iii) For the Bessel initial state, the oscillation mode appears, and the amplitude can be modulated by the strength of the imaginary transverse field and the localization length of the initial state. Our results reveal the mechanism of magnetic BOs in the non-Hermitian quantum spin chain and pave the way for future research on BOs in other quantum systems.

### v2

@@ III.3 Bessel initial state
While the results in Eqs. (27) and (28) indicate that this is a periodic oscillation with period $T=\pi/h_{z}$, the pattern is not so explicit. Nevertheless, for a finite system size $N$, we introduce the center of the wave packet:
@@ IV Numerical simulations
The boundary of $\sigma^{z}_{j}\right>=+1$ and $-1$ is the position of the magnetic domain wall. For the Hermitian spin chain, Figs. 4(a1)-(a3) show that the dynamics are magnetic Bloch breathing, BO, and stationary modes for the Kronecker delta, Gaussian, and Bessel initial states, respectively. With the same initial states, for the non-Hermitian spin chain, the results in Figs. 4(b1)-(b3) indicate that the dynamics are magnetic Bloch breathing, stationary, and BO modes, respectively. For the latter two initial states, the corresponding BOs exhibit totally different patterns for the same initial states in the two different systems. These numerical results are in accordance with the analyses in the previous section.
Figure 5: Overlap of perturbative solutions and numerically evolved states at instant $t=10$ for three different initial states considered in the main text. (a) Overlap for the Hermitian system as a function of real transverse field $g$. (b) Overlap for the non-Hermitian system as a function of imaginary transverse field $g$. Parameters of initial states are taken as the same as those in Fig. 4. Other parameters of the system are set as $N=16$, $J=1$, and $h_{z}=0.2$.
In order to further estimate the validity of perturbative solutions presented in the previous section, we compare the numerical evolved states with perturbative solutions by the overlap
$O=\left|\langle\Psi_{\textrm{num.}}\left(t\right)|\Psi_{\textrm{ana.}}\left(t\right)\rangle\right|$ (31)
at instant $t$, where $|\Psi_{\textrm{num.}}\left(t\right)\rangle$ and $|\Psi_{\textrm{ana.}}\left(t\right)\rangle$ denote the normalized numerical and analytical evolved states, respectively. The perturbative solutions $|\Psi_{\textrm{ana.}}\left(t\right)\rangle$ for three initial states are taken as the form of Eq. (16) with $f_{m}(t)$ being Eqs. (III.1), (23) and (27), respectively. In Fig. 5, we presented the numerical results of overlap $O$ at instant $t=10$ as a function of real and imaginary $g$. It indicates that for both cases, the perturbative solutions are in good agreement with the exact numerical results in a small $\left|g\right|$, while for imaginary $g$ the overlap drops sharply when $\left|g\right|\gtrsim h_{z}/2=0.1$ due to the $\mathcal{PT}$ symmetry breaking of $H$ that is not captured by the effective Hamiltonian $H_{\textrm{eff}}$.
@@ V Summary
In summary, we demonstrate the existence of the magnetic BOs in a non-Hermitian quantum Ising chain. It is shown that in the small-field region, the low-energy dynamics of the magnetic domain walls are captured by a single-particle effective Hamiltonian, with the transverse field acting as a hopping coefficient for the magnetic domain wall and the strength of the longitudinal field playing the role of a skew potential. For real and imaginary transverse fields, the eigenstates of the effective Hamiltonian are both localized states with equally spaced energy levels, forming the WS ladders. Analytical and numerical calculations of the time evolution for the non-Hermitian quantum Ising chain show the following.
(i) The dynamics of the Kronecker delta initial state follow a breathing mode.
(ii) The Gaussian state remains stationary, which is different from the oscillation of the domain wall in a Hermitian quantum Ising chain.
(iii) For the Bessel initial state, the oscillation mode appears, and the amplitude can be modulated by the strength of the imaginary transverse field and the localization length of the initial state.
The validity of perturbative solutions is estimated by comparing them with numerical results. Our results reveal the mechanism of magnetic BOs in the non-Hermitian quantum spin chain and pave the way for future research on BOs in other quantum systems.